OpenStudy (anonymous):

factor if possible: c^6 -343

OpenStudy (anonymous):

here, c^6-343=c^6-7^3

OpenStudy (anonymous):

Well, 343 is 7 cubed. So we have a difference of cubes as c^6 = (c^2)^3. A difference of cubes is factorised as follows: a^3-b^3 = (a-b)(a^2+ab+b^2) So, applying above, c^6-7^3 = (c^2-7)(c^4+7c^2+49)

OpenStudy (anonymous):

(c^2)^3- 7^3

OpenStudy (anonymous):

a^3-b^3 = (a-b)(a^2+ab+b^2

OpenStudy (anonymous):

whoa!! i'm lost...

OpenStudy (anonymous):

a= c^2 , b = 7

OpenStudy (anonymous):

plug ,a= c^2 , b = 7 in a^3-b^3 = (a-b)(a^2+ab+b^2

OpenStudy (anonymous):

we have c^6-343=(c^2-7)(c^4+7c^2+49)

OpenStudy (anonymous):

got it now.?